The bearing of a lighthouse from ship A is N29°E and the bearing of that lighthouse from ship B is N61°W. If ship A and ship B are on the east-west line 200 kilometers apart, how far is ship B from the lighthouse?


1
Expert's answer
2021-10-26T02:28:37-0400


Let the distance be denoted by X which is the Adjacent side of the triangle.

Use the formula; "SOH\ CAH\ TOA, \\"

"Cos\theta=\frac{Adjecent}{Hypotenuse} \implies Cos\ 29^{\circ}\ = \frac{Adjacent}{200}\\
\implies200\ Cos\ 29^{\circ}=Adjacent\\
\therefore\ Adjacent=174.92km"

Below is an image to show the triangle.